OpenStudy (anonymous):

I have a problem that says, as x goes toward infinity, the function approaches (?) The rational function is f(x)= 4x^2-1/4x^2-9 How do I figure out how to answer that!? Help please! =(

OpenStudy (anonymous):

then you have juste to calculate 4x^2/4x^2 =1

terenzreignz (terenzreignz):

Here's how to do it... \[\Large \frac{4x^2-1}{4x^2 - 9}\]

OpenStudy (anonymous):

can you tell me why? I don't really get it if i don't know why.

terenzreignz (terenzreignz):

Figure out the largest exponent of x, whether in the numerator or the denominator. .. so... what IS the largest exponent of x? ^_^

OpenStudy (anonymous):

2

terenzreignz (terenzreignz):

That's right. So what I want you to do now, is multiply that entire thingy with this *ugly* expression... \[\Huge \frac{\frac1{x^2}}{\frac1{x^2}}\]

terenzreignz (terenzreignz):

By the way, that *ugly* expression up there^ is effectively equal to 1

OpenStudy (anonymous):

so just multiply it by 1?

terenzreignz (terenzreignz):

Yeah... but a rather special incarnation of 1. You'll see why later...

OpenStudy (agent0smith):

You can also do it by just ignoring everything but the highest powers of x on the numerator and denominator (these become the dominant terms as x gets larger) in this case it's... \[\Large \frac{ 4x^2 }{ 4x^2 }= \]

terenzreignz (terenzreignz):

\[\huge \frac{4x^2 - 1}{4x^2 - 9}\cdot \frac{\frac1{x^2}}{\frac1{x^2}}\] And this is just equal to \[\Huge \frac{\frac{4x^2}{x^2}-\frac1{x^2}}{\frac{4x^2}{x^2}-\frac9{x^2}}\] Which just looks horrible doesn't it? :P

OpenStudy (anonymous):

i'm not very good at multiplying fractions and things. i'm pathetic really. failing college algebra for the 6th time.

terenzreignz (terenzreignz):

But do you see what I did there? I just multiplied everything (in the numerator and denominator) by \(\Huge \frac1{x^2}\)

OpenStudy (agent0smith):

Then do my method @kawriter

terenzreignz (terenzreignz):

Yeah, I suppose ^ I do love the magic, though :3

OpenStudy (anonymous):

i'm not sure I see how you got what you got.

OpenStudy (anonymous):

is that the only step @agent0smith ?

OpenStudy (agent0smith):

Look at the highest powers of x on the top and bottom. Ignore everything else.

OpenStudy (agent0smith):

You have to simplify what i gave.

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

The thing is, @kawriter as x gets larger and larger, those 'lesser' exponents become more and more insignificant, and that the value of the expression depends more on the x with the highest exponent...so you really only have to check how that 'highest-exponent' term behaves ^_^

OpenStudy (anonymous):

@agent0smith doI simplify it by square rooting the whole thing and coming out with 2/2 = 1?

OpenStudy (anonymous):

crap! now i have to graph that function... If i'm asking too much of anyone, just let me know!

OpenStudy (agent0smith):

Yes. Another example: \[\Large y = \frac{ -2x^3 + 3x^2-x+1 }{ 4x^4-7x^3-5 }\] First identify the highest power on the top and bottom.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

and 4

OpenStudy (agent0smith):

\[\Large y = \frac{ -2x^3 + 3x^2-x+1 }{ 4x^4-7x^3-5 }\] highest power of x on the top is the -2x^3. What about the bottom?

OpenStudy (anonymous):

4^4

OpenStudy (anonymous):

that graphing page is AMAZING!! THANK YOU!

OpenStudy (agent0smith):

You mean 4x^4?

OpenStudy (anonymous):

yes sorry.

OpenStudy (agent0smith):

Now divide them like i did above, for the first question.

OpenStudy (anonymous):

so would that be -1/2x?

OpenStudy (agent0smith):

Yep. And what is -1/2x close to if x is huge?

OpenStudy (anonymous):

um, i have no idea.....3?

OpenStudy (agent0smith):

When you divide a small number by a huge number, what's the result close to?

OpenStudy (anonymous):

the huge number

OpenStudy (anonymous):

okay, i think.

OpenStudy (agent0smith):

So a small number divided by a huge number is close to...?

OpenStudy (anonymous):

an even smaller number?

OpenStudy (agent0smith):

Yes, but you could round it to what?

OpenStudy (anonymous):

1?

OpenStudy (agent0smith):

Well is a number like 0.000000001 closer to 0 or 1?

OpenStudy (anonymous):

oh, 0

OpenStudy (agent0smith):

Which means -1/2x, if x is huge, is basically equal to...

OpenStudy (anonymous):

0

OpenStudy (agent0smith):

Yep. What about 10/2x^2, if x is huge, it must be basically...

OpenStudy (anonymous):

0

OpenStudy (anonymous):

so for my initial problem, the function approaches 0?

OpenStudy (agent0smith):

You worked out the answer to the original one earlier... "doI simplify it by square rooting the whole thing and coming out with 2/2 = 1?"

OpenStudy (agent0smith):

It's only equal to zero if AFTER simplifying, you end up with an x on the bottom, like -10/3x.

OpenStudy (anonymous):

gotcha. i just did another one and I got 10/x so that equals 0?

OpenStudy (agent0smith):

Yes. And if you end up with something like x/5... what result do you get if you divide a huge number by a small number?

OpenStudy (anonymous):

a really big number? cause x isn't on the bottom.

OpenStudy (agent0smith):

Yes, which just means the function approaches infinity as x approaches infinity.

OpenStudy (anonymous):

coooooollllll!!! makes sense!

OpenStudy (anonymous):

and the one i did earlier,

OpenStudy (anonymous):

i had 4x^2/4x^2

OpenStudy (anonymous):

so i took the square root of both, x^2's cancelled out and I was left with 2/2 which =1

OpenStudy (agent0smith):

You can't take the square root of anything, all you can do is cancel off common factors. The 4s cancel, and the x^2 both cancel, leaving you with 1.

OpenStudy (anonymous):

ohhhhhhhh.....gotcha. okay. why isn't it 0 if both cancel?

OpenStudy (agent0smith):

Well... is 4 divided by 4 equal to zero?

OpenStudy (anonymous):

oh. you make sense sensei. lol. i'm super dumb in math! sorry!

OpenStudy (agent0smith):

Remember that x is just a number. So is x^2. 20x^5/20x^5 = 1

OpenStudy (anonymous):

ahhh....cool.

OpenStudy (anonymous):

thanks @agent0smith you are AWESOME!

OpenStudy (agent0smith):

You're welcome :)

OpenStudy (anonymous):

can i ask you a new question?

OpenStudy (agent0smith):

Sure.

OpenStudy (anonymous):

is the remainder theorem like synthetic division?

OpenStudy (anonymous):

if it's not i'm so screwed. ha.

OpenStudy (agent0smith):

Well, you can use synthetic division to get a result/remainder, if that's what you mean.

OpenStudy (anonymous):

okay, i think that's what i mean. heh. i'm given an equation and i have to use the theorem to find f(-3)

OpenStudy (agent0smith):

Yeah, you can use synth division. The remainder will be the value of f(-3)

OpenStudy (anonymous):

oh okay cool.

OpenStudy (anonymous):

starting synth division always messes me up. do you start out multiplying?

OpenStudy (agent0smith):

Like this: http://www.purplemath.com/modules/polysolv/remndr02.gif The first coefficient, you just bring it down, then multiply by the number you're dividing by.

OpenStudy (agent0smith):

I got that here: http://www.purplemath.com/modules/remaindr.htm

OpenStudy (anonymous):

would you mind checking it?

OpenStudy (anonymous):

i'm so bookmarking that page! thanks

OpenStudy (agent0smith):

It has several examples of synthetic division at the bottom. What was the actual question?

OpenStudy (anonymous):

Given 9x^4 + 12x^3 +2x^2 -7x +74, find f(-3)

OpenStudy (anonymous):

no, I got 292 =(

OpenStudy (anonymous):

Ahhh!!! I don't know what I did wrong! =(

OpenStudy (agent0smith):

You could take a pic of your work and upload it on photobucket or something, and paste the link here so i can check.

OpenStudy (anonymous):

my phone can't take pictures. =(

OpenStudy (anonymous):

Can I maybe describe my work?

OpenStudy (anonymous):

i had the -15 as -39. wrong from the start. okay i'm going to work it one more time without looking at the picture.

OpenStudy (agent0smith):

Fixed an error